OpenStudy (anonymous):

The vertex is 2,9 and the points are -2,5. and you have to convert it to y-k=a(x-h)^2

OpenStudy (bahrom7893):

okay do u remember what are h and k?

OpenStudy (anonymous):

the points right

OpenStudy (bahrom7893):

well no they're coordinates of a point. Which point, do u remember that?

OpenStudy (anonymous):

okay i get that the points that i put up there are h,k and i just want to see if i got the problem right because my teacher doesnt want us to have fractions soo yea

OpenStudy (bahrom7893):

okay h and k are coordinates of the vertex, so: h = 2, k = 9 a is the slope to find the slope from two points (2,9) and (-2,5): Slope = rise/run = (y-y1)/(x-x1) = (5 - 9)/(-2-2) = -4/-4=1

OpenStudy (bahrom7893):

y-k=a(x-h)^2 y-9=1(x-2)^2, or y-9=(x-2)^2